The freight train headed south at 9 a.m. and the express train headed north from the same station at noon. At 3 p.m., the trains were 420 miles apart. What was the speed of each if the speed of the express train was 20 mph greater than the speed of the freight train?
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

The freight train headed south at 9 a.m. and the express train headed north from the same station at noon. At 3 p.m., the trains were 420 miles apart. What was the speed of each if the speed of the express train was 20 mph greater than the speed of the freight train?
Let speed of freight train be x mph
speed of express train = (x+20 ) mph
after 6 hours they were 420 miles away from each other
Distance = speed * time
They are movig away from each other.
6x+6(x+20)=420
6x+6x+120=420
12x =300
speed of freight train = 25 mph
speed of express train = 25+20 =45 mph



Answer by greenestamps(13206)



Another wrong answer from tutor @mananth, who provides responses without checking to see if her answers are correct....

(The freight train travels for 6 hours, but the express train only travel for 3 hours.)

6(x)+3(x+20) = 420
6x+3x+60 = 420
9x = 360
x = 40

ANSWERS:
freight train: x = 40 mph
express train: x+20 = 60 mph

CHECK: 6(40)+3(60) = 240+180 = 420